OpenStudy (anonymous):

Calculus 2 Help please!!!

OpenStudy (anonymous):

\[\int\limits\frac{ \sin^5(x) }{1+\cos^2(x) }dx\]

OpenStudy (anonymous):

I got this far: \[\int\limits \frac{ (1-\cos^2)^2 }{ 1+\cos^2 }dx\] u=cos(x) du=-sin(x)dx dx=-1/sin(x)du \[\int\limits \frac{ (1-u^2)^2 }{1+u^2 }du\] I don't know what to do after this...

OpenStudy (anonymous):

That isn't quite right. \[\int\limits \frac {\sin^5 (x)}{1 + \cos^2 (x)}dx = \int\limits \frac {(1- \cos^2 (x))^2 \sin (x)}{1+\cos^2(x)}dx\]The rest of your problem looks good.

OpenStudy (anonymous):

XD thats what I meant to type

OpenStudy (anonymous):

Also, make sure you specify that \[dx = -\frac {du}{sin(x)}\]

OpenStudy (anonymous):

From here, try and use integration by parts to solve this equation.

OpenStudy (anonymous):

Also, need to add a negative sign to your equation.

OpenStudy (anonymous):

so would it just turn the whole integration negative?

OpenStudy (anonymous):

Yes. This is because \[dx = - \frac {du}{sin(x)}\]the negative sign on this substitution.

OpenStudy (anonymous):

So.. \[-\int\limits\limits \frac{ 1-2u^2+u^4 }{ 1+u }du\] =\[-\int\limits(\frac{ 1 }{ 1+u^2 })du +\int\limits(\frac{ 2u^2 }{ 1+u^2})du-\int\limits \frac{ u^4 }{ 1+u^2 }du\]

OpenStudy (anonymous):

first one is integration of tan inverse x...I am stuck with the rest 2...

OpenStudy (anonymous):

With \[\int\limits \frac{ 2u^2 }{ 1+u^2 }du\] would we split it like\[\int\limits \frac{ 2 }{ 1+u^2 }\frac{ u^2 }{ 1+u^2 } du\]

OpenStudy (anonymous):

no can't do that...

OpenStudy (anonymous):

That isn't right. That is the equivalent of \[\int \frac {2u^2}{(1+u^2)^2}\]What you need to do is use integration by parts: \[\int u(x)v'(x) = u(x)v(x) - \int u'(x)v(x)dx\]

OpenStudy (anonymous):

how do you guys use the integration symbols in here teach me quick...

OpenStudy (anonymous):

what would we use as u and v

OpenStudy (anonymous):

lol got it..

OpenStudy (anonymous):

You can either use the Equation button beneath the chat box to input equations, or use LaTeX notation to manually type it out.

OpenStudy (anonymous):

\[u(x) = 2u^2\] or the numerator and \[v(x) = \frac {1}{1+u^2}\]

OpenStudy (anonymous):

\[\left(\frac{u^2-\left(1+u^2\right)}{1+u^2}+1\right)\]

OpenStudy (anonymous):

\[\left(2\int\limits \frac{u^2-\left(1+u^2\right)}{1+u^2}du+\int\limits \:1du\right)\]

OpenStudy (anonymous):

wow its so hard to type these stuff

OpenStudy (anonymous):

It's nice once you get used to it. I'm off for the night. Good luck. I know the solution, but I didn't do the work to get it, if you want me to post it.

OpenStudy (anonymous):

form that you get tan inverse u + x

OpenStudy (anonymous):

please give me the solution

OpenStudy (anonymous):

\[\int\limits2u^2/(1+u2)\]

OpenStudy (anonymous):

thats how u break that

OpenStudy (anonymous):

\[\int \frac {sin^5(x)}{1+cos^2(x)} = \frac {cos^3(x)}{3} -3cos(x) +4tan^{-1}(cos(x)) + C\]Courtesy of Wolfram Alpha

OpenStudy (anonymous):

Thank you for helping, Good night

OpenStudy (anonymous):

Another solution by plugging the original equation into Wolfram Alpha is \[\int \frac {sin^5(x)}{1+cos^2(x)}dx = \frac {11cos(x)}{4} - \frac {1}{12}cos(3x) - 4tan^{-1}(cos(x)) +C\]